OpenStudy (kanwal32):

The sum of first n terms of the series 1^2+2.2^2+3^2+2.4^2.........is n(n+1)^2, when n is even. When n is odd

OpenStudy (kanwal32):

the sum is

OpenStudy (kanwal32):

@Queen_Bee1 this is the question

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

well i got 20.6

OpenStudy (kanwal32):

no but we have to represent the answer in terms of n

OpenStudy (kanwal32):

@wio help

OpenStudy (anonymous):

What exactly is the sequence?

OpenStudy (kanwal32):

\[1^{2}+2.2^{2} + 3^{2}+2.4^{2}.......\]

OpenStudy (anonymous):

2.2?

OpenStudy (anonymous):

What is the sequence in terms of \(n\)?

OpenStudy (kanwal32):

sum is n(n+1)^2/2 when n is even we have to fnd when n is odd

OpenStudy (anonymous):

Ok, if \(n\) is odd, then it is just \(S_{n+1} - a_n\)

OpenStudy (kanwal32):

ok

OpenStudy (anonymous):

And \(S_{n+1}\) is even, so you can use your formula.

OpenStudy (anonymous):

When \(n\) is odd, it seems \(a_n\) is \(n^2\).

OpenStudy (anonymous):

Does that help you?

OpenStudy (kanwal32):

can u pls write the equation

OpenStudy (anonymous):

You want a direct answer?

OpenStudy (kanwal32):

yes

OpenStudy (anonymous):

Okay, plug \(n+1\) into the answer for the even sum. Then subtract \(n^2\). That is the answer.

OpenStudy (kanwal32):

pls can u write direct answer

OpenStudy (anonymous):

I'm too lazy.

OpenStudy (kanwal32):

pls

OpenStudy (anonymous):

\[ \frac{(n+1)((n+1)+1)^2}2 - n^2\]

OpenStudy (kanwal32):

but the answer is \[n^2(n+1)/2\]

OpenStudy (anonymous):

wait a moment, is that the answer for even or odd?

OpenStudy (kanwal32):

for odd

OpenStudy (anonymous):

hmmm, well mine needs to be simplified.

OpenStudy (anonymous):

that or the even formula was wrong?

OpenStudy (kanwal32):

is that giving the same

OpenStudy (anonymous):

oh wait. I think I should have done \(-a_{n+1}\)

OpenStudy (anonymous):

so \(-2n^2\)

OpenStudy (anonymous):

Eh, I didn't put a lot of thought into it.

OpenStudy (kanwal32):

@ganeshie8 PLS HELP

ganeshie8 (ganeshie8):

i think wio's method is using triangular numbers, and after representing the nth term you need to find the sum...

OpenStudy (anonymous):

My method is finding the sum for \(n+1\), since we know the even sum, and then just subtracting the last \(n+1\) term, to get the \(n\) sum.

ganeshie8 (ganeshie8):

Oh okay... looks i misinterpreted the solution

OpenStudy (anonymous):

\[ S_n = S_{n+1}-a_{n+1} \]

OpenStudy (anonymous):

If \(n\) is odd then\(n+1\) is even. We know how to do the even sum. That was my logic.

ganeshie8 (ganeshie8):

wow! yes last term is tricky, got you :)

OpenStudy (kanwal32):

got u @wio

OpenStudy (kanwal32):

this question Is a bit tricky and I have a test ON 6TH

OpenStudy (kanwal32):

can any 1 write the solution

OpenStudy (kanwal32):

pls

OpenStudy (kanwal32):

@dan815 @hartnn @ganeshie8 @ParthKohli @batmann

ganeshie8 (ganeshie8):

did u get wio's logic ?

OpenStudy (kanwal32):

yes

Parth (parthkohli):

is this the series?\[1^2 + (2\times 2^2) + 3^2 + (2\times 4^2)\cdots\]

OpenStudy (kanwal32):

no I have written thee series above

ganeshie8 (ganeshie8):

You're given a formula which works when \(n\) is even : \[\large 1^2+2.2^2+3^2+2.4^2 + \cdots + 2.n^2= n(n+1)^2 \]

OpenStudy (kanwal32):

yes

OpenStudy (kanwal32):

its

ganeshie8 (ganeshie8):

and you're trying to find a formula that works when \(n\) is odd : \[\large 1^2+2.2^2+3^2+2.4^2 + \cdots + 2.(n-1)^2 + n^2 = ? \]

OpenStudy (anonymous):

write it like this (1^2+2^2+3^2+...+n^2) + (2^2+4^2+6^2+...n^2) Now, try to find the sum.

OpenStudy (kanwal32):

yes

Parth (parthkohli):

\[= (1^2 + 2^2 + 3^2 + \cdots n^2) + (2^2 + 4^2 + 6^2\cdots n ^2)\]

ganeshie8 (ganeshie8):

we're already given the formula for sum when "n" is even

OpenStudy (kanwal32):

yeah

ganeshie8 (ganeshie8):

we just need to use it and deduce a formula for odd "n"

OpenStudy (kanwal32):

ok

ganeshie8 (ganeshie8):

which was exactly what @wio did above ^

OpenStudy (kanwal32):

thnx @ganeshie8 @wio

OpenStudy (ikram002p):

if you have it like this \(\sum_{n=1}^n n(n+1)^2 \) why you need to find odd /even ?

ganeshie8 (ganeshie8):

let me rewrite the wio's solution again in one reply

OpenStudy (ikram002p):

or this is what your qn asking for ? how to find sum when n is odd and find the sum when n is even ?

ganeshie8 (ganeshie8):

\[\large \color{red}{1^2+2.2^2+3^2+2.4^2 + \cdots + 2.(n-1)^2} + n^2 \] \[\large \color{red}{\dfrac{(n-1)(n-1+1)^2}{2}}+ n^2 \] \[\large \color{red}{\dfrac{(n-1)(n)^2}{2}}+ n^2 \] \[\large \color{red}{\dfrac{(n-1)(n)^2 + 2n^2}{2}}\] \[\large \color{red}{\dfrac{n^2(n+1)}{2}}\]

ganeshie8 (ganeshie8):

ikram : we're given a formula for `even n`, our job is to deduce the formula for `odd n`

OpenStudy (kanwal32):

thnx for u support every1

OpenStudy (ikram002p):

oh , ic ^^ for once get confused

ganeshie8 (ganeshie8):

lol same wid me, i went straight for the proof before wio hinting me

OpenStudy (ikram002p):

huh

OpenStudy (kanwal32):

ok

OpenStudy (kanwal32):

should I close the problem